OpenStudy (anonymous):

a colony of bacteria grows exponentially according to the function N(t)=100e^0.045t where N measures in grams and t is measured in days. 1. determine the initial amount of bacteria?? 2. what is the population after 5 days?? 3. how long will it take the population to reach 140 grams??

OpenStudy (anonymous):

1. The initial amount would be when t=0. N(0)=100e^0 = 100*1 = 100 grams of bacteria initially. 2. t=5 \[N(5)=100e^{.045*5}\] N(5)=100*1.252323 N(5)=125.2323 grams of bacteria. 3. N(t)=140 140=100e^.045t 1.4=e^.045t t is approximately 1.3384, so 1.3384 days.